OpenStudy (anonymous):

Transform the formula for the given variable: a = (1/2)bh Solve for h. A. h = (2a)/b B. h = 2a + b C. h = 2a – b D. h = a/b

OpenStudy (anonymous):

A

OpenStudy (anonymous):

a = (1/2)bh a\b=(1/2)h 2a\b=h

OpenStudy (anonymous):

Thanks(: